OpenStudy (anonymous):

limx-> 1 (1/x-3) + (1/x^2-3x+2)

OpenStudy (anonymous):

\[ \dfrac{1}{x-3} + \dfrac{1}{x^2-3x+2}=-\dfrac{1}{x-1}+\dfrac{1}{x-2}+\dfrac{1}{x-3}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

give me the medal :3

OpenStudy (anonymous):

lol it is infinity

OpenStudy (anonymous):

the second term has a zero in the denominator, and it doesn't go away when you add it to the first term, so the limit is infinite

OpenStudy (anonymous):

yeah \[\lim_{x -> 1}\dfrac{1}{x-1}=\infty\]

OpenStudy (anonymous):

hey what about my medal?

OpenStudy (anonymous):

@darthjavier \to

OpenStudy (anonymous):

\[\lim_{x\to 1}f(x)\]

OpenStudy (anonymous):

\[\text{OKAY}\]

OpenStudy (anonymous):

also \rigtharrow i think, but \to is shorter